OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2). y = 5/2 x - 2 y = -5/2 x + 8 Y = 2/5 x - 6 y = -2/5 x + 4

OpenStudy (anonymous):

slope intercept form is y=mx+c

OpenStudy (anonymous):

where m is slope and y is intercept on y axis

OpenStudy (anonymous):

u know slope?

OpenStudy (anonymous):

y=mx+b*

OpenStudy (anonymous):

Determine your slope \[\frac{ y2-y1 }{ x2-x1 } = slope\]

OpenStudy (anonymous):

so m=-2-8/0-4

OpenStudy (anonymous):

m=10/4 m=5/2

OpenStudy (anonymous):

Sorry im just writing stuff down lol

OpenStudy (anonymous):

so eq is y=5/2x-2